# FARM: Few-shot Adaptive Malware Family Classification under Concept Drift

### v1

@@ Abstract
Malware classification models often face performance degradation due to concept drift, arising from evolving threat landscapes and the emergence of novel malware families. This paper presents FARM (Few-shot Adaptive Recognition of Malware), a framework designed to detect and adapt to both covariate and label drift in Windows Portable Executable (PE) malware classification. FARM leverages a triplet autoencoder to project samples into a discriminative latent space, enabling unsupervised drift detection via DBSCAN clustering and dynamic thresholding. For rapid adaptation, it employs few-shot learning using prototype-based classification, requiring only a handful of labeled samples. FARM also supports full retraining when enough drifted samples accumulate, updating the latent space for long-term integration. Experiments on the BenchMFC dataset demonstrate that FARM improves classification performance under covariate drift by 5.6%, and achieves an average F1 score of 0.85 on unseen malware families using only few-shot adaptation, which further increases to 0.94 after retraining. These results highlight FARM’s robustness and adaptability in dynamic malware detection environments under limited supervision.
@@ I Introduction
The malware domain is inherently dynamic. Malware continuously evolves, either through the emergence of entirely new malware families with unknown characteristics or through the modification of existing malware behaviors by malware authors [47]. If malware detection and family attribution were based on a static view of the domain, these tasks might not appear particularly challenging. Indeed, leveraging machine learning (ML) and deep learning (DL) techniques, malware can often be detected with high accuracy when evaluated on static datasets. However, when faced with the real-world scenario of a constantly evolving malware landscape, the reliability and performance of malware detectors decrease significantly [15].
Empirical studies underscore the severity of this issue. For instance, when a malware classifier is trained on data collected at an initial time point and tested on future data, its detection performance deteriorates substantially, by up to 50% compared to traditional randomized train-test splits that ignore temporal context [35]. Similarly, another notable study indicates that within just six months of deployment, the performance of a malware detection model dropped from 99% to 76% [7]. Such findings clearly illustrate the significant impact that evolving malware characteristics have on model effectiveness, emphasizing the necessity to address malware detection as a continuously adapting task.
@@
Label drift describes the emergence of entirely new, previously unknown malware families, introducing novel patterns that detection models have never encountered. This form of drift challenges classifiers by presenting entirely new classes requiring models to adapt or retrain to maintain their predictive capabilities. Covariate drift, on the other hand, refers to evolving behaviors and properties of existing malware families, where malware authors alter characteristics to evade detection while preserving the overall malware class identity [21, 17]. Both types of drift necessitate ongoing adaptation or retraining of malware classifiers to sustain high detection accuracy and to effectively mitigate evolving cybersecurity threats.
@@
The primary challenges in mitigating concept drift and enhancing malware detection involve both model adaptation and data labeling. One key issue is updating the model to reflect evolving patterns while accurately determining whether a malware sample is novel or drifted. Relying solely on probability scores from machine learning or deep learning models is problematic, as these scores, especially from softmax-based classification layers, tend to remain high and overconfident even for previously unseen samples [22, 34, 18]. Additionally, determining the optimal frequency for model retraining is challenging; too frequent updates may result in insufficient data for effective learning, while infrequent updates may fail to capture significant changes in malware behavior [22]. Adding to these difficulties, labeling samples is expensive and time-consuming, particularly given the large datasets required by conventional deep learning approaches [23]. Moreover, labels provided by antivirus vendors or human experts are often unreliable, inconsistent, or prone to errors [27].
Our approach addresses these challenges by initially focusing on the detection of drifted samples through metric learning, specifically using triplet networks. By computing the distance between a sample and established cluster centers, the model can discern whether a sample belongs to a known malware family or should be classified as unknown or drifted, thereby mitigating the issue of misleading high probability scores. Furthermore, to reduce the burden of extensive data labeling and to enable agile adaptation to evolving malware behaviors, our method integrates meta-learning via prototypical learning. This strategy allows the model to adapt effectively with only a few new samples, ensuring both efficiency and responsiveness in dynamic environments.
@@
While the majority of prior work on concept drift and adaptation in malware detection has focused on the Android platform, our study targets Windows Portable Executable (PE) malware. The daily detection rate of Windows PE malware samples exceeds 300,000, based on industry-reported data [4]. These new samples encompass both variants of existing malware families, exhibiting novel behavioral patterns or characteristics, and completely novel malware families previously unseen by detection models. Both scenarios pose substantial challenges for existing malware classifiers, negatively impacting their generalization capabilities and reliability in real-world deployment. To the best of our knowledge, this is the first comprehensive framework that combines drift detection and adaptation tailored specifically to Windows PE malware, addressing a significant gap in the existing literature. In summary, our work makes the following contributions:
@@
  • We propose Few-shot Adaptive Recognition of Malware (FARM), the first comprehensive framework that combines drift detection and adaptation tailored specifically to Windows PE malware.
  • We develop an approach for detecting previously unseen malware samples by leveraging triplet deep neural networks, which accurately identify both drifted and novel instances.
  • We propose an efficient model update mechanism that requires only a few new samples, thereby enhancing the model’s performance on emerging malware families.
  • We provided an integrated adaptation framework that combines efficient few-shot adaptation with long-term retraining.
  • We validate our methodology on a timestamped and meticulously labeled dataset that captures drifted malware, demonstrating robust performance in dynamic, real-world scenarios.
@@ II-C Few-shot Adaptation
While prior works have explored concept drift in malware detection, much of this research has centered on the Android platform [22, 5, 47, 7, 17]. Existing approaches have focused on detecting drift but often lack rapid adaptability or fine-grained class modeling [47, 7]. Few-shot learning has been explored in cybersecurity, but rarely in ways that directly target the evolving structure of Windows PE malware. FARM fills this gap by providing a comprehensive approach to drift detection and adaptation for Windows PE malware, effectively handling both covariate and label drift. FARM leverages metric learning for unsupervised drift detection and incorporates few-shot learning to rapidly adapt to new or evolving malware families using only a small number of labeled samples. By combining clustering with dynamic thresholding, FARM achieves a strong balance between rapid adaptation and stable performance, demonstrating robust generalization under minimal supervision on the BenchMFC [21] dataset.
@@ III Methodology
Our approach involves multiple components, primarily the training of a triplet autoencoder network, which is utilized for both classification and concept drift detection, alongside a prototypical network to adapt the model to new data distributions. An overview of the FARM pipeline is illustrated in Figure 1. The system takes EMBER features [2] as input, which are then embedded into a latent space using a triplet-trained encoder. In the Drift Detection phase, DBSCAN is used to identify dense clusters and detect outliers or shifted samples. If a sample lies outside known cluster thresholds, it is flagged as drifted and passed to the Drift Adaptation module. There, drifted samples are stored in a buffer and periodically clustered to discover coherent groups. Once a stable cluster is detected, a prototype is generated to represent a new or evolved malware family, allowing the classifier to integrate the new behavior.
@@ III-A Training phase
$L(x_{a},x_{p},x_{n})=\max\left\{0,\,D(x_{a},x_{p})-D(x_{a},x_{n})+\alpha\right\}$
@@
$L_{\text{triplet}}=\frac{1}{N}\sum_{i=1}^{N}L(x_{a_{i}},x_{p_{i}},x_{n_{i}})$
@@
$L_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\|x_{i}-\hat{x}_{i}\|^{2}$
The combined loss function $L_{Final}$, incorporating both triplet and reconstruction objectives, guides the network to produce embeddings that are both compact for classification and informative for reconstruction [14].
$L_{Final}=L_{triplet}+\lambda L_{\text{MSE}}$
@@ III-B Drift detection phase
Rather than relying on traditional clustering techniques such as K-means, which assumes spherical clusters of roughly equal size and requires the number of clusters to be specified a priority, we employ the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [11]. DBSCAN is particularly well-suited to this task due to its ability to identify clusters of arbitrary shapes and densities without requiring the number of clusters in advance. This is especially important in our setting, where a single malware family may exhibit internal diversity, leading to multiple natural groupings within the same class. In addition, DBSCAN’s ability to distinguish between core samples, border points, and noise makes it effective for detecting outliers and potentially novel behaviors. Moreover, DBSCAN has demonstrated strong performance in malware clustering tasks [12, 25], further reinforcing its suitability for our application. This clustering process is visualized in Figure 2, where samples are embedded using a triplet autoencoder and grouped into clusters via DBSCAN.
Figure 2: Illustration of the clustering and thresholding process within FARM. After projection into the latent space by the triplet autoencoder, DBSCAN identifies dense regions representing clusters. For each cluster, a centroid (red cross) and a maximum intra-cluster distance are computed, forming a threshold boundary for classification and outlier rejection.
@@
$\tau_{i}=\max_{x\in C_{i}}\|f(x)-\mu_{i}\|^{2}$
This threshold $\tau_{i}$ effectively defines the radius of the smallest hypersphere that encompasses all members of cluster $C_{i}$, excluding outliers identified by DBSCAN. During inference, a new sample $x_{\text{new}}$ is embedded via $f(x_{\text{new}})$ and assigned to the closest cluster $C_{i}$ only if:
$\|f(x_{\text{new}})-\mu_{i}\|^{2}\leq\tau_{i}$
Otherwise, it is rejected as an outlier or flagged for further analysis. This strategy enables a principled, data-driven approach to defining class boundaries and integrates naturally with the metric space learned by the triplet autoencoder.
@@ III-C Drift adaptation phase
The drift adaptation phase addresses the emergence of novel malware families not represented in the initial training distribution. Instead of retraining the model upon the arrival of each new class, the system adopts a two-stage strategy that balances rapid response with long-term integration. Initially, EMBER features extracted from a PE malware are embedded into a latent space using the triplet encoder. If the sample is identified as drifted, meaning it exceeds the threshold of all known clusters, it is stored in a buffer for further processing. Upon each addition to the buffer, DBSCAN clustering is re-applied to the accumulated drifted samples. Unlike the clustering performed during the initial drift detection phase, this instance of DBSCAN is used to discover cohesive groupings among the buffered samples, specifically targeting the identification of latent clusters containing at least 10 samples. Once such a cluster is found, its mean embedding is computed to form a new prototype, which is then added to the set of known clusters. This enables future samples to be classified into the newly identified group, thereby allowing the model to adapt to novel malware families without full retraining. To distinguish between representations of known and novel families, we refer to the known cluster means as centroids and to the novel representations as prototypes. Although both are computed as class-wise means, this terminological distinction reflects their respective roles within the system.
@@
As described previously, when a cohesive cluster of drifted samples is detected, a prototype is formed and incorporated into the set of known malware family representations. This prototype is then used during inference to classify incoming samples. The system continuously monitors these assignments and accumulates samples that are classified into the novel family. Once the number of such samples reaches a threshold sufficient to support stable retraining, thereby avoiding overfitting or catastrophic forgetting, the model undergoes a full retraining procedure.
@@ IV Evaluation and Experimental Setup
  • RQ1: How effectively can metric learning, particularly through a triplet autoencoder, detect concept drift in malware classification? This question is addresses in Section V.
@@
  • RQ3: To what extent can FARM generalize to previously unseen malware behaviors and maintain robust performance in real-world deployment settings? This question is addressed in Section VI-C
@@ IV-A Dataset Description
For our experiments, we utilize a subset of the BenchMFC dataset [21], which comprises over 223,000 unpacked malware samples spanning 526 known families, collected over the period from 2012 to 2022. Each sample in the dataset is annotated with its corresponding malware family and a timestamp indicating its earliest known appearance, enabling temporal segmentation of malware evolution. BenchMFC covers a broad range of malware types and families, supporting evaluations of classification systems under distributional changes.
In our setup, we focus exclusively on unpacked samples and do not consider packing-related drift or packed variants, as packing can significantly impact concept drift and has been shown to challenge the generalization ability of static machine learning-based detectors [1]. We therefore leave the exploration of such effects to future work. In line with the setup described by [21], we adopt the same controlled subset of the dataset, which includes 16 malware families, as summarized in Table I. The dataset consists of three categories:
  • one representing known families used during training,
  • one representing evolving families used during covariate drift evaluation,
  • and one representing unseen families used during label drift evaluation.
Notably, both the malware family labels and the drift categorization are adopted directly from the original BenchMFC dataset, with no additional labeling applied. Each category includes eight malware families, with 500 Windows PE malware samples per family, forming a balanced dataset suitable for evaluating metric learning and adaptation strategies under both covariate and label drift scenarios.
@@
yuner 500 yuner 500 vofbus 500

### v2

@@ Abstract
Malware classification models often suffer performance degradation under concept drift due to evolving threat landscapes and the emergence of novel malware families. This paper presents FARM (Few-shot Adaptive Recognition of Malware), a unified framework for detecting and adapting to both covariate drift and label drift in Windows Portable Executable (PE) malware family classification. FARM uses a triplet autoencoder to project samples into a discriminative latent space, enabling unsupervised drift detection through DBSCAN clustering and dynamic thresholding. To enable rapid adaptation, the framework employs a few-shot strategy that can incorporate new classes from only a small number of labeled samples. FARM also supports full retraining when sufficient drifted samples accumulate, allowing longer-term model updating. Experiments on the BenchMFC dataset show that FARM improves classification performance under covariate drift by 5.6%, and achieves an average F1 score of 0.85 on unseen malware families using few-shot adaptation, increasing to 0.94 after retraining. These results indicate that FARM provides an effective approach for drift-aware malware family classification in dynamic environments with limited supervision.
@@ I Introduction
The malware domain is inherently dynamic. Malware continuously evolves, either through the emergence of entirely new malware families with unknown characteristics or through the modification of existing malware behaviors by malware authors [47]. If malware detection and family attribution were based on a static view of the domain, these tasks might not appear particularly challenging. Indeed, leveraging machine learning and deep learning techniques, malware can often be detected with high accuracy when evaluated on static datasets. However, when faced with the real-world scenario of a constantly evolving malware landscape, the reliability and performance of malware detectors decrease significantly [15, 22, 35].
Empirical studies underscore the severity of this issue. For instance, when a malware classifier is trained on data collected at an initial time point and tested on future data, its detection performance deteriorates substantially, by up to 50% compared to traditional randomized train-test splits that ignore temporal context [35]. Similarly, another notable study indicates that within just six months of deployment, the performance of a malware detection model dropped from 99% to 76% [7]. Such findings clearly illustrate the significant impact that evolving malware characteristics have on model effectiveness, emphasizing the necessity to address malware detection as a drift-aware adaptation problem, where models must detect distributional change and update their decision structure over time.
@@
Label drift describes the emergence of entirely new, previously unknown malware families, introducing novel patterns that detection models have never encountered. This form of drift challenges classifiers by presenting entirely new classes that require models to adapt or retrain to maintain their predictive capabilities. Covariate drift, on the other hand, refers to evolving behaviors and properties of existing malware families, where malware authors alter characteristics to evade detection while preserving the overall malware class identity [21, 17]. Both types of drift necessitate ongoing adaptation or retraining of malware classifiers to sustain high detection accuracy and to effectively mitigate evolving cybersecurity threats.
@@
The primary challenges in mitigating concept drift and enhancing malware family detection involve both model adaptation and data labeling. One key issue is updating the model to reflect evolving patterns while accurately determining whether a malware sample is novel or drifted. Relying solely on probability scores from machine learning or deep learning models is problematic, as these scores, especially from softmax-based classification layers, tend to remain high and overconfident even for previously unseen samples [22, 34, 18]. Additionally, determining the optimal frequency for model retraining is challenging; too frequent updates may result in insufficient data for effective learning, while infrequent updates may fail to capture significant changes in malware behavior [22]. Adding to these difficulties, labeling samples is expensive and time-consuming, particularly given the large datasets required by conventional deep learning approaches [23]. Moreover, labels provided by antivirus vendors or human experts are often unreliable, inconsistent, or prone to errors [27].
Our approach addresses these challenges by first detecting potentially drifted samples in a learned metric space, using distances to family-level cluster representations rather than relying solely on classifier confidence. Samples that fall outside the accepted region of known families are treated as candidates for further analysis rather than being forcefully assigned to an existing class. To support rapid response under limited supervision, the framework then introduces a few-shot adaptation mechanism that can represent emerging malware behaviors using only a small number of labeled samples. When sufficient evidence accumulates, the model is retrained to consolidate these additions into the latent space. Accordingly, FARM is designed as a drift-aware adaptation pipeline with both immediate-response and longer-term consolidation stages.
@@
While the majority of the prior work on concept drift and adaptation in malware detection has focused on the Android platform [22, 5, 47, 7, 17], our study focuses on Windows Portable Executable (PE) malware. According to industry-reported statistics, the daily detection volume of Windows PE malware exceeds 300,000 samples [4]. These incoming samples may correspond either to variants of existing malware families, which exhibit new behavioral patterns or characteristics, or to previously unseen malware families. Both cases can degrade the performance of static malware classifiers and reduce their reliability in real-world deployment.
Motivated by this challenge, we propose a unified framework for detecting and adapting to concept drift in Windows PE malware classification. Rather than treating drift detection and model adaptation as separate problems, our framework integrates both components to support rapid response to emerging samples and longer-term model updating. In summary, our work makes the following contributions:
@@
  • We propose Few-shot Adaptive Recognition of Malware (FARM), a framework for Windows PE malware family classification under concept drift.
  • We combine latent-space drift detection with few-shot adaptation to support the recognition of previously unseen malware families and evolving family variants.
  • We evaluate the proposed framework on a timestamped malware benchmark under covariate-drift and label-drift scenarios, demonstrating its effectiveness in temporally evolving classification settings.
@@ II-C Few-shot Adaptation
While prior works have explored concept drift in malware detection, much of this research has focused on the Android platform [22, 5, 47, 7, 17]. Existing approaches have primarily emphasized drift detection or continual updating, often without explicitly coupling drift identification with rapid adaptation and fine-grained malware family modeling [47, 7]. Few-shot learning has also been studied in cybersecurity, but has been less explored in the context of temporally evolving Windows PE malware family classification.
In this context, FARM is designed as a unified framework for Windows PE malware family classification under concept drift. It combines metric-learning-based drift detection with few-shot adaptation to support rapid incorporation of new or evolving malware families using only a small number of labeled samples. In addition, FARM integrates clustering and dynamic thresholding for drift identification together with a longer-term retraining strategy, enabling the framework to address both covariate drift and label drift within a single pipeline. We evaluate this design on the BenchMFC [21] dataset to study its effectiveness in temporally evolving malware classification settings.
@@ III Methodology
Our approach involves multiple components, primarily the training of a triplet autoencoder network, which is utilized for both classification and concept drift detection, alongside a few-shot strategy to adapt the model to new data distributions. An overview of the FARM pipeline is illustrated in Figure 1. The system takes EMBER features [2] as input, which are then embedded into a latent space using a triplet-trained encoder. In the drift detection phase, DBSCAN is used to identify dense clusters and detect outliers or shifted samples. If a sample lies outside known cluster thresholds, it is flagged as drifted and passed to the drift adaptation module. There, drifted samples are stored in a buffer and periodically clustered to discover coherent groups. Once a stable cluster is detected, a prototype is generated to represent a new or evolved malware family, allowing the classifier to integrate the new behavior.
@@ III-A Training phase
$L(x_{a},x_{p},x_{n})=\max\left\{0,\,D(x_{a},x_{p})-D(x_{a},x_{n})+\alpha\right\}$ (1)
@@
$L_{\text{triplet}}=\frac{1}{N}\sum_{i=1}^{N}L(x_{a_{i}},x_{p_{i}},x_{n_{i}})$ (3)
@@
$L_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\|x_{i}-\hat{x}_{i}\|^{2}$ (4)
The combined loss function $L_{Final}$, incorporating both triplet and reconstruction objectives, guides the network to produce embeddings that are discriminative for classification while remaining informative enough for reconstruction [14].
$L_{Final}=L_{triplet}+\lambda L_{\text{MSE}}$ (5)
@@ III-B Drift detection phase
Rather than relying on traditional clustering techniques such as K-means, which require the number of clusters to be specified a priori, we employ the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [11]. The primary motivation for using DBSCAN in our framework is its ability to identify an unknown number of dense subclusters directly from the embedding space, without requiring a fixed subgroup count for each malware family. This is particularly important in our setting, where a single malware family may exhibit multiple dense subclusters in the latent space due to internal behavioral diversity. Furthermore, DBSCAN distinguishes between core samples, border points, and noise, which is useful for reducing the influence of sparse or atypical samples during cluster construction. Prior studies have also reported strong performance of DBSCAN in malware clustering contexts [12, 25], further supporting its use in our framework. Accordingly, DBSCAN is adopted here as a practical mechanism for discovering family-level substructure under unknown complexity in the latent space. This clustering process is illustrated in Figure 2, where samples are embedded using a triplet autoencoder and then grouped via DBSCAN.
Figure 2: DBSCAN-derived subclusters and centroid-based acceptance thresholds in FARM.
@@
$\tau_{i}=\max_{x\in C_{i}}\|f(x)-\mu_{i}\|^{2}$ (8)
This threshold $\tau_{i}$ defines a cluster-specific acceptance boundary around the centroid of $C_{i}$, determined by the maximum distance of its non-outlier members in the embedding space. During inference, a new sample $x_{\text{new}}$ is mapped to $f(x_{\text{new}})$ and assigned to the cluster with the nearest centroid only if:
$\|f(x_{\text{new}})-\mu_{i}\|^{2}\leq\tau_{i}$ (9)
Otherwise, it is rejected as an outlier and flagged for further analysis. This strategy provides a lightweight, data-driven approximation for cluster assignment that is consistent with the metric space learned by the triplet autoencoder.
@@ III-C Drift adaptation phase
The drift adaptation phase addresses the emergence of novel malware families not represented in the initial training distribution. Instead of retraining the model upon the arrival of each new class, the system adopts a two-stage strategy that balances rapid response with long-term integration. This design targets drift-aware adaptation under evolving malware behavior, where rapid response and later consolidation are both important, but it is not intended to constitute a full classical continual learning framework with long sequential task updates and explicit forgetting-aware optimization. Initially, EMBER features extracted from a PE malware are embedded into a latent space using the triplet encoder. If the sample is identified as drifted, meaning it exceeds the threshold of all known clusters, it is stored in a buffer for further processing. Upon each addition to the buffer, DBSCAN clustering is re-applied to the accumulated drifted samples. Unlike the clustering used during the initial drift detection phase, DBSCAN is applied here to discover cohesive groupings within the buffered samples. The minimum cluster size is set to 10, which serves as a practical compromise between robustness and early adaptation: smaller groups are more likely to be unstable, while larger requirements would delay recognition of emerging malware families. This choice is consistent with the few-shot analysis in Section VI-C1, where performance improves with additional support samples, making 10 a reasonable balance between accuracy and responsiveness. Once such a cluster is found, its mean embedding is computed to form a new prototype, which is then added to the set of known clusters. This enables future samples to be classified into the newly identified group, thereby allowing the model to adapt to novel malware families without full retraining. To distinguish between representations of known and novel families, we refer to the known cluster means as centroids and to the novel representations as prototypes. Although both are computed as class-wise means, this terminological distinction reflects their respective roles within the system.
@@
As described previously, when a cohesive cluster of drifted samples is detected, a prototype is formed and incorporated into the set of known malware family representations. This prototype is then used during inference to classify incoming samples. The system continuously monitors these assignments and accumulates samples that are classified into the novel family. Once the number of such samples reaches a threshold sufficient to support stable retraining, the model undergoes a full retraining procedure.
@@ IV Evaluation and Experimental Setup
  • RQ1: How effectively can metric learning, particularly through a triplet autoencoder, detect concept drift in malware classification? This question is addressed in Section V.
@@
  • RQ3: To what extent can FARM adapt to previously unseen malware families under label drift using only a small number of labeled samples? This question is addressed in Section VI-C
@@ IV-A Dataset Description
For our experiments, we use the BenchMFC dataset [21] as the source benchmark. BenchMFC comprises over 223,000 unpacked malware samples spanning 526 known families, collected between 2012 and 2022. Each sample is annotated with its malware family and a timestamp indicating its earliest known appearance, enabling temporal segmentation of malware evolution. The dataset covers a broad range of malware types and families, making it suitable for evaluating classification systems under distributional change.
In our setup, we focus exclusively on unpacked samples and do not consider packing-related drift or packed variants, as packing can significantly impact concept drift and has been shown to challenge the generalization ability of static machine learning-based detectors [1]. We leave the exploration of such effects to future work. Following [21], we adopt the same controlled 16-family subset summarized in Table I. This subset is organized into three categories:
  • known families used for training,
  • evolving families used for covariate drift evaluation, and
  • unseen families used for label drift evaluation.
Both the malware family labels and the drift categorization are adopted directly from the original BenchMFC dataset, with no additional labeling applied. Each category contains eight malware families, with 500 Windows PE malware samples per family, forming a balanced dataset for evaluating metric learning and adaptation strategies under both covariate and label drift scenarios.
@@
yuner 500 yuner 500 vobfus 500